OpenStudy (anonymous):

Help please! anti derivative of sinx/cos^2/3 x

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{\sin x}{(\cos x)^{2/3}}dx\quad=\quad \int\limits \frac{1}{(\color{royalblue}{\cos x})^{2/3}}\left(\color{orangered}{\sin x\;dx}\right)\] We need to make a u-sub, \[\Large\bf\sf \color{royalblue}{u=\cos x}\]

zepdrix (zepdrix):

What part are you stuck on? :)

OpenStudy (anonymous):

I get stuck on the 2/3

zepdrix (zepdrix):

Were you able to get your substitution setup correctly? Do you know how to find your du?

OpenStudy (anonymous):

du is -sinxdx

zepdrix (zepdrix):

Ok good, moving the negative over, \[\Large\bf\sf \color{orangered}{-du=\sin x\;dx}\]

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits \frac{1}{(\color{royalblue}{\cos x})^{2/3}}\left(\color{orangered}{\sin x\;dx}\right)\quad=\quad \int\limits\limits \frac{1}{(\color{royalblue}{u})^{2/3}}\left(\color{orangered}{-du}\right)\]

zepdrix (zepdrix):

So the 2/3 is messing you up? Just bring it up to the numerator using rules of exponents.\[\Large\bf\sf \frac{1}{x^n}\quad=\quad x^{-n}\]From there you can simply apply the power rule!! :)

OpenStudy (anonymous):

I get 3/5 u^5/3

zepdrix (zepdrix):

Hmm did you remember that your u is negative? :O I don't think you'll get all the way up to 5/3 power. \[\Large\bf\sf \int\limits u^{-2/3}\;du\quad=\quad \frac{1}{-\frac{2}{3}+1}u^{-2/3+1}\quad=\quad \frac{1}{1/3}u^{1/3}\] \[\Large\bf\sf =\quad 3u^{1/3}\]

zepdrix (zepdrix):

Your exponent is negative* i meant :p

OpenStudy (anonymous):

Duh! I forgot the negative! okay.. so is 3cos^1/3 x the antiderivative?

zepdrix (zepdrix):

Mmm ya looks good! With a +c somewhere at the end.

OpenStudy (anonymous):

Thank You!